Movies, Inc. is considering a new project producing educational movies for children. The project requires an investment in equipment of $1,800,000. The equipment will be fully depreciated using straight line depreciation over its 8 year accounting life. At the end of the project the firm expects to be able to sell the equipment for $125,000. The firm expects to sell 475,000 million videos each year for 5 years. The retail price for each video will be $20.00 and variable costs related to production are $12.00 per video. Additional fixed costs include $900,000 per year for labor, $450,000 for rent and $200,000 for all other expenses. The project also requires an additional investment in net working capital of $700,000, all of which will be returned to the firm at the end of the project. The firm’s marginal tax rate is 34% and the required return on this project is 12%, compounded quarterly. Would you recommend this project to the firm? Why or why not? Clearly explain your answer.
Answer #1
Present Value(PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=discount rate=Required Return =12%=0.12
N=Year of cash flow
a Initial Outlay
A Investment in equipment $1,800,000
B Investment in Net working Capital $700,000
C=A+B Initial Outlay Associated with the project $2,500,000
Cash Flow in Year 1-4
a Annual Sales quantity(units)               475,000,000
b Contribution per unit=20-12 $8
c=a*b Total Contribution margin $3,800,000,000
d Additional fixed cost for labor -$900,000
e Rent expense -$450,000
f Other fixed expenses -$200,000
g Annual Depreciation=(1800000)/8 -$225,000
h=c+d+e+f+g Annual Before tax Income $3,798,225,000
i=h*(1-0.34) Annual after tax Income $2,506,828,500
j Add: Depreciation (Non cash expense) $225,000
k=i+j After Tax Cash Flow $2,507,053,500
Terminal Cash Flow in Year5: $0
l Annual after tax cash flow $2,507,053,500
m Release of net working capital $700,000
n Salvage value of equipment $125,000
o Accumulated depreciation=225000*5 $1,125,000
p Book Value at end of 5 years $675,000 (1800000-1125000)
q=p-n Loss on salvage $550,000
r=q*34% Tax saving on salvage $187,000
s=n+r Terminal Cash flow on salvage $312,000
t=l+m+s Terminal Cash Flow in Year5: $2,508,065,500
N Year 0                              1                              2                              3                               4                              5
CF Net Cash Flow -$2,500,000 $2,507,053,500 $2,507,053,500 $2,507,053,500 $2,507,053,500 $2,508,065,500 SUM
PV=CF/(1.12^N) Present Valure -$2,500,000 $2,238,440,625 $1,998,607,701 $1,784,471,162 $1,593,277,823 $1,423,143,721 $9,035,441,031
NPV=Sum of PV Net Present Value $9,035,441,031
d Project NPV $9,035,441,031
NPV is POSITIVE
Yes, this project is recommended